# Optimal Stopping for Sequential Bayesian Experimental Design

### v1

@@ 5.2 Contaminant Source Detection in Convection-Diffusion Field
Figure 4: Training convergence of the convection-diffusion problem. Top: average reward; bottom: average stopping stage. Columns correspond to different horizons NN and experimental costs.
For zero cost ($c_{k}=0$), Figure 4 (left column) shows that both vanilla and curriculum learning converge successfully. The vanilla PG approach quickly settles onto the maximum stage of $N=4$, while curriculum PG achieves slightly higher rewards. In this regime, continuation rewards dominate termination rewards, even under poor initial designs, so both methods are naturally incentivized to continue and eventually converge.
For negative costs, the training challenges identified in our theoretical analysis become evident. With moderate cost ($c_{k}=-0.5$, middle column), vanilla PG fails to outperform a fixed stage-4 stopping policy and prematurely convergence to stage 3. In contrast, curriculum learning sustains exploration longer, achieving higher rewards and more appropriate stopping behavior that balances information gain with cost. At higher cost ($c_{k}=-0.8$, right column), vanilla method degrades further, stopping too early, whereas curriculum learning maintains stable convergence and superior reward.
These results show that sequential dependence and cost-sensitive stopping can destabilize joint optimization, trapping vanilla PG in suboptimal policies. Curriculum learning overcomes this by gradually relaxing stopping constraints, enabling robust convergence from exploration to exploitation.
@@ 6 Conclusion
This work tackles the fundamental challenge of deciding when to terminate sequential experiments in Bayesian experimental design. We developed a principled framework for optimal stopping that accounts for the expected future value of continuation, showing that the optimal rule is simple: continue only when the expected continuation reward exceeds the immediate terminal reward.
From a computational perspective, the coupling between design and stopping policies creates circular dependencies that destabilize training. We addressed this with a PG approach augmented with curriculum learning, which gradually transitions from forced continuation to adaptive stopping. Our experiments demonstrate that, especially for tasks with strong sequential dependencies, curriculum learning achieves stable convergence. The convection-diffusion source detection problem illustrates this setting, where curriculum learning yields substantial gains over fixed stopping rules by breaking the coupling during training.
While our framework offers a principled solution, several limitations point to future work. First, evaluating terminal rewards at every decision stage can be expensive with costly posterior updates or high-dimensional reward structures. Second, alternative formulations that directly parameterize the stopping policy without relying on the optimal stopping theorem (Theorem 1) might avoid circular dependencies and simplify training. Such approaches may converge more slowly or sacrifice theoretical optimality, but could broaden the range of practical applications—an avenue that remains open for investigation.
Overall, this framework advances autonomous experimental systems by enabling intelligent, resource-aware stopping decisions, improving efficiency of sequential Bayesian experimental design.
@@ Appendix A Appendix
A.1 Proof of Theorem 1
Proof of Theorem 1.
@@ A.1 Proof of Theorem 1
By induction, $V^{\pi,\psi}_{k}(s_{k})\geq V^{\pi,\psi^{\prime}}_{k}(s_{k})$ for all $k=0,\ldots,N-1$. In particular, $V^{\pi,\psi}_{0}(s_{0})\geq V^{\pi,\psi^{\prime}}_{0}(s_{0})$, which proves that the stopping policy in Theorem 1 is optimal. ∎
A.2 Proof of Theorem 2
Proof of Theorem 2.
We first prove the stopping sets are equivalent under two formulations. For convenience, we omit the notation of design policy $\pi$ and stopping policy $\psi$, but distinguish them using upperscript of incr and term respectively. The policy dependence in the expectations is also omitted for notational consistency. We first prove the following relationship between value functions:
@@ A.2 Proof of Theorem 2
$V^{incr}_{k+1}(s_{k+1})=V^{term}_{k+1}(s_{k+1})-\left[D_{\mathrm{KL}}(p_{\theta|I_{k+1}}\,||\,p_{\theta|I_{0}}))+\sum_{i=0}^{k}c_{i}(\xi_{i})\right].$ (32)
@@
Therefore, the stopping sets are equivalent in both formulations. Especially, the optimization objective is
@@
A.3 Proof of Theorem 3
To avoid notation congestion, below we will omit the subscript on $w$ and shorten $\mu_{k,w_{k}}(s_{k})$ to $\mu_{k,w}(s_{k})$, with the understanding that $w$ takes the same subscript as the $\mu$ function. We also omit explicit conditioning in expectations to save formula space, with the understanding that expectations are taken with respect to the appropriate policies and states.
Proof of Theorem 3.
The gradient of the expected utility can be written using the V-function:
@@ A.3 Proof of Theorem 3
$V_{k}^{\pi_{w},\psi_{w}}(s_{k})$ $=\bm{1}_{s_{k}\in\mathcal{T}_{k,w}}r^{S}_{T}(s_{k})+(1-\bm{1}_{s_{k}\in\mathcal{T}_{k,w}})\mathbb{E}_{y_{k}}\left[r^{C}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k})+V^{\pi_{w},\psi_{w}}_{k+1}(\mathcal{F}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k}))\right],$ (39)
the gradient of $V_{k}^{\pi_{w},\psi_{w}}(s_{k})$ is
$\quad\nabla_{w}V^{\pi_{w},\psi_{w}}_{k}(s_{k})$ $=\nabla_{w}\bm{1}_{s_{k}\in\mathcal{T}_{k,w}}r^{S}_{T}(s_{k})-\nabla_{w}\bm{1}_{s_{k}\in\mathcal{T}_{k,w}}\mathbb{E}_{y_{k}}\left[r^{C}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k})+V^{\pi_{w},\psi_{w}}_{k+1}(\mathcal{F}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k}))\right]$
$\quad+(1-\bm{1}_{s_{k}\in\mathcal{T}_{k,w}})\nabla_{w}\mathbb{E}_{y_{k}}\left[r^{C}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k})+V^{\pi_{w},\psi_{w}}_{k+1}(\mathcal{F}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k}))\right]$
$=\nabla_{w}\bm{1}_{s_{k}\in\mathcal{T}_{k,w}}\left(r^{S}_{T}(s_{k})-\mathbb{E}_{y_{k}}\left[r^{C}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k})+V^{\pi_{w},\psi_{w}}_{k+1}(\mathcal{F}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k}))\right]\right)$
$\quad+\bm{1}_{s_{k}\notin\mathcal{T}_{k,w}}\nabla_{w}\mathbb{E}_{y_{k}}\left[r^{C}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k})+V^{\pi_{w},\psi_{w}}_{k+1}(\mathcal{F}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k}))\right].$ (40)
@@
$h(s_{k},w)=r^{S}_{T}(s_{k})-\mathbb{E}_{y_{k}}\left[r^{C}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k})+V^{\pi_{w},\psi_{w}}_{k+1}(\mathcal{F}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k}))\right]=0.$ (41)
The indicator function changes value only on this boundary. Therefore, the gradient can be expressed as a Dirac delta function centered on the boundary:
$\nabla_{w}\bm{1}_{s_{k}\in\mathcal{T}_{k,w}}=\delta(h(s_{k},w))\nabla_{w}h(s_{k},w).$ (42)
The first term in (40) is then reduced to $h(s_{k},w)\delta(h(s_{k},w))\nabla_{w}h(s_{k},w)$, which is always 0 with the property of the Dirac Delta function $x\delta(x)\equiv 0$. For the second term in (40),
$\quad\nabla_{w}\mathbb{E}_{y_{k}}\left[r^{C}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k})+V^{\pi_{w},\psi_{w}}_{k+1}(\mathcal{F}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k}))\right]$
$=\nabla_{w}\Bigg[\int_{y_{k}}p(y_{k}|s_{k},\mu_{k,w}(s_{k}))r^{C}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k})\,dy_{k}+\int_{s_{k+1}}p(s_{k+1}|s_{k},\mu_{k,w}(s_{k}))V^{\pi_{w},\psi_{w}}_{k+1}(s_{k+1})\,ds_{k+1}\Bigg]$
$=\nabla_{w}\int_{y_{k}}p(y_{k}|s_{k},\mu_{k,w}(s_{k}))r^{C}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k})\,dy_{k}+\nabla_{w}\int_{s_{k+1}}p(s_{k+1}|s_{k},\mu_{k,w}(s_{k}))V^{\pi_{w},\psi_{w}}_{k+1}(s_{k+1})\,ds_{k+1}$

### v2

@@ 4.2 One-dimensional nonlinear test problem
Figure 6: Training convergence of the one-dimensional nonlinear test problem. Top: average reward; bottom: average stopping stage. Columns correspond to different horizons NN and experimental costs.
Figure˜6 summarizes three representative settings. With zero experimental cost and horizon $N=3$, both vanilla PG and curriculum PG rapidly reach comparable rewards, and the learned stopping stage remains close to the terminal horizon. This is consistent with the nonnegative information-gain reward: in the absence of experimental cost, additional measurements are beneficial unless the posterior has already become sufficiently concentrated.
When a constant cost $c_{k}=-0.5$ is imposed for the same horizon, the optimal behavior shifts toward early termination. Both methods converge to stopping after approximately one experiment, indicating that the first nonlinear observation already provides enough information to make further measurements unattractive relative to their cost. Curriculum PG initially explores longer trajectories, as intended, and then transitions to the same early-stopping regime once the stopping rule is fully activated.
For the longer horizon $N=5$ with moderate cost $c_{k}=-0.2$, both methods converge to similar final rewards and stopping stages, terminating around stage two. However, the training curves exhibit larger fluctuations. This behavior is consistent with a flatter reward landscape, where stopping at adjacent stages can yield nearly indistinguishable total rewards, analogous to the linear-Gaussian case with $N=4$ and $c_{k}=-0.25$. In the nonlinear problem, this trade-off is further complicated by sequential dependence, since each observation changes the posterior shape and hence the value of subsequent designs. As a result, small policy changes can shift trajectories among several nearly equivalent stopping behaviors, producing larger fluctuations even though both methods ultimately settle to comparable performance.
4.3 Contaminant Source Detection in a Convection-Diffusion Field
We next consider a sensor-movement problem for contaminant source detection, which exhibits stronger sequential dependence than the preceding examples. In the linear-Gaussian benchmark, the optimal design and stopping decisions are independent of the realized observations. In the one-dimensional nonlinear problem, observations affect future decisions through the updated posterior. Here, the dependence is stronger: the design variable is a sensor displacement, so the current sensor location is determined by the cumulative history of all previous designs, in addition to the information contained in past observations. Mobile sensors must therefore be strategically repositioned to localize an unknown pollution source, with each placement decision depending critically on both previous measurements and previously selected movements.
Contaminant transport is governed by a convection-diffusion PDE on a two-dimensional square domain $z_{L},z_{R}]^{2}$:
$\frac{\partial G(z,t;\theta)}{\partial t}=\nabla^{2}G-u(t)\cdot\nabla G+S(z,t;\theta),\qquad\qquad z\in[z_{L},z_{R}]^{2},\quad t>0$ (28)
where $\theta=[\theta_{x},\theta_{y},\theta_{h},\theta_{s}]\in\mathbb{R}^{4}$ parameterizes a Gaussian source
$S(z,t;\theta)=\frac{\theta_{s}}{2\pi\theta_{h}^{2}}\exp\!\left(-\frac{(\theta_{x}-z_{x})^{2}+(\theta_{y}-z_{y})^{2}}{2\theta_{h}^{2}}\right)$ (29)
with location $(\theta_{x},\theta_{y})$, width $\theta_{h}$, and strength $\theta_{s}$, and $u=[u_{x},u_{y}]\in\mathbb{R}^{2}$ is a time-dependent convection velocity. A mobile sensor measures noisy concentrations
$y_{k}=G(z=s_{k+1,p},\,t_{k};\,\theta)+\epsilon_{k},\qquad\epsilon_{k}\sim\mathcal{N}(0,\sigma_{\epsilon}^{2}),$ (30)
where the design variable $\xi_{k}$ represents sensor displacement,
$s_{k+1,p}=s_{k,p}+\xi_{k}.$ (31)
We consider $N$ measurement opportunities at times $t_{k}=k\cdot\Delta t$, $k=0,\ldots,N-1$, with $\Delta t=5.0\times 10^{-2}$, and fix $\theta_{h}=0.05$, $\theta_{s}=2$, $u_{x}=u_{y}=10t/0.2$, with prior $\theta_{x},\theta_{y}\sim\mathcal{U}(0,1)$.
The forward model (28) is solved numerically using a second-order finite volume method on a uniform grid with spatial resolution $dz_{x}=dz_{y}=0.01$, together with a second-order fractional-step method for time integration using a timestep of $dt=5.0\times 10^{-4}$. Since repeated evaluations of the forward model are computationally expensive, we construct deep neural network (DNN) surrogate models for $G(z,t_{k};\theta)$ at each stage $k$. Each surrogate takes a 4-dimensional input $(z,\theta)$ and maps it through five hidden layers containing 40, 80, 40, 20, and 10 neurons, respectively, to produce a scalar concentration output. The training dataset is generated from 2000 prior samples of $\theta$, restricted to the reachable domain and split 80%/20% for training and testing. Figure˜7 compares concentration contours at four time points obtained from the finite volume solver (upper panel) and the DNN surrogate (lower panel). The two sets of contours are visually nearly indistinguishable, with test mean-squared errors on the order of $10^{-6}$. The surrogate models also provide a computational speedup of approximately $10^{5}$ relative to the finite volume solver.
Figure 7: Comparison of the concentration field G at 4 time points with $\theta_{x}=0.43$ and $\theta_{y}=0.22$, obtained using the finite volume (upper panel) and DNN surrogate (lower panel). The true solution is solved using finite volume in the wider computational domain $1,2]^{2}$ but displayed here in the region of interest $0,1]^{2}$.
Figure 8: Training convergence of the convection-diffusion problem with horizon $N=4$. Top: average reward; bottom: average stopping stage. Columns correspond to different experimental costs.
For zero cost ($c_{k}=0$), Figure 8 (left column) shows that both vanilla and curriculum learning converge successfully to the terminal horizon. Vanilla PG reaches the high-reward regime rapidly, while curriculum PG initially follows the forced-continuation schedule and then approaches a comparable reward. In this regime, continuation rewards dominate termination rewards, so even imperfect early designs provide sufficient information to favor continuing through the full horizon.
For negative costs, the training challenges identified in Section 3.2 become more evident. These cases reveal reward landscapes with multiple local optima associated with different stopping stages. The design variables are sensor movements rather than independent sensor locations, so each early movement changes the feasible locations and measurements available at later stages. If the critic underestimates the value of continuation early in training, the induced stopping rule truncates trajectories before the policy experiences informative later-stage moves. Those missing trajectories then reinforce the underestimated continuation value, creating an early-stopping local optimum.
With moderate cost ($c_{k}=-0.5$, middle column), the two local optima are clearly separated in reward. Vanilla PG converges to the earlier, lower-reward optimum, stopping at an average stage of approximately 2.4. In contrast, curriculum PG forces longer trajectories early in training, allowing the policy and critic to discover a later-stopping optimum with average stopping stage slightly above 3 and higher reward. This is the clearest case where the circular dependency produces a practical failure mode: vanilla PG becomes trapped in the smaller early-stopping optimum and misses the higher-reward solution discovered by curriculum PG.
At higher cost ($c_{k}=-0.8$, right column), the learning dynamics again indicate two local optima, as vanilla PG and curriculum PG converge to different average stopping stages. In this case, however, the two optima have similar objective values, so the final rewards are comparable. Vanilla PG reaches the earlier-stopping optimum, while curriculum PG converges to the later-stopping optimum because its training trajectories are forced to continue longer before the stopping rule is fully activated. Thus, curriculum learning helps avoid missing the global or higher-reward optimum by preserving long-horizon exploration before early stopping decisions are allowed to dominate training. In both negative-cost cases, the learned policies outperform fixed-stage stopping rules, indicating that the optimal stopping strategy in this problem is state-dependent rather than associated with a single fixed stopping stage.
(a) Vanilla, $\theta=(0.61,0.73)$, total reward = -0.564 (information gain = 0.436, experimental cost = -1.0)
(b) Curriculum, $\theta=(0.61,0.73)$, total reward = 2.198 (information gain = 3.698, experimental cost = -1.5)
Figure 9: Episode instances for vanilla (upper) and curriculum (lower) policies in the $c_{k}=-0.5$ scenario. Purple star: true source location; red dots: sensor positions; red lines: sensor movements; contours: posterior PDF.
Figure˜9 presents representative episodes generated using the policies learned by the vanilla PG and curriculum PG methods for the case $c_{k}=-0.5$. Starting from the same initial condition, the two methods produce slightly different experimental designs at the initial stage, indicating that different stopping-learning strategies lead to different learned design policies. This further supports the circular dependency discussed earlier: premature stopping affects the learning of the design policy itself. The vanilla PG policy terminates after two experiments with a total reward of -0.564, while the curriculum-trained policy continues for three experiments and achieves a higher total reward of $2.198$.
We further explore a more complex cost structure in the contaminant source detection problem with design-dependent experimental costs $c_{k}=-3\|\xi_{k}\|$, where longer sensor movements incur larger penalties. Figure˜10 shows training convergence under this cost structure. Both vanilla and curriculum PG converge to similar rewards and stopping stages, with the average stopping stage approaching approximately 3.6–3.7. The design-dependent cost also changes the learned movement strategy. Figure˜11 compares the spatial distribution of learned sensor movements under constant versus design-dependent costs across experimental stages. At stage $k=0$, the fixed initial state makes the learned movement deterministic within each cost setting, although the selected initial movement differs between the two cost structures. Under the constant cost (upper panels), later-stage movements become more dispersed, especially for $k=2$ and $k=3$, since movement length is not directly penalized. In contrast, under the design-dependent cost (lower panels), the learned movements concentrate along a narrow trajectory with much smaller magnitudes, reflecting a policy that balances information acquisition against the linear movement penalty. This adaptive behavior demonstrates that our framework successfully incorporates spatially varying costs into the joint design and stopping optimization.
These results demonstrate that the proposed framework can accommodate complex, design-dependent cost structures. They also show that the benefit of curriculum learning depends on the interaction between cost, stopping, and sequential design dependence: in this design-dependent-cost setting, vanilla and curriculum training converge to similar stopping stages and rewards, while in regimes where premature stopping suppresses later-stage exploration, curriculum learning can provide a clear advantage.
Figure 10: Training convergence of the convection-diffusion source detection problem with design-dependent costs $c_{k}=-3\|\xi_{k}\|$. Left: average reward; right: average stopping stage.
Figure 11: Comparison of learned sensor movement policies under constant versus design-dependent costs in the convection-diffusion source detection problem. Upper panels: constant cost $c_{k}=-0.5$. Lower panels: design-dependent cost $c_{k}=-3\|\xi_{k}\|$.
Overall, these results demonstrate that strong sequential dependence combined with cost-sensitive stopping can destabilize joint optimization, trapping vanilla PG in suboptimal early-stopping policies. Curriculum learning mitigates this effect by preserving longer training trajectories before adaptive stopping dominates, with the clearest gains appearing when premature stopping would otherwise suppress later-stage exploration. To illustrate the role of the curriculum schedule, we compared several common choices for the stopping probability $p_{\text{stop}}(\ell)$ in Figure˜12, including linear, exponential, and sigmoid increases. As shown in Figure˜13, these schedules behave similarly on relatively simple problems such as the linear-Gaussian benchmark: all converge to the optimal policy, differing mainly in convergence patterns and speed. However, for more complex or long-horizon tasks such as the contaminant source detection problem, the choice of schedule becomes critical. As shown in Figure˜14, a fast exponential schedule leads to insufficient exploration during early training, resulting in training instability and convergence to a suboptimal policy. In contrast, sigmoid and linear schedules with slower progression rates ensure adequate full-horizon exploration in early stages, leading to stable training and superior final performance. These results suggest that practitioners should tune the curriculum schedule based on problem complexity: simple problems tolerate aggressive schedules, while complex tasks benefit from gradual transitions that maintain exploration before allowing the optimal stopping rule to dominate training dynamics.
Figure 12: Examples of stopping probability schedules for curriculum PG over 300 training iterations.
Figure 13: Comparison of different curriculum schedules on the Linear-Gaussian benchmark. Left: average reward; right: average stopping stage.
Figure 14: Comparison of different curriculum schedules on the convection-diffusion source detection problem. Left: average reward; right: average stopping stage.
@@ 5 Conclusion
This paper addressed the problem of deciding when to terminate a sequential Bayesian experimental design campaign. The central idea is to treat stopping as part of the sequential decision problem rather than as a fixed threshold rule. This leads to an interpretable optimal stopping criterion: experimentation should continue only when the expected continuation value exceeds the immediate terminal reward.
The resulting optimization problem is challenging because the design policy, continuation value, and stopping boundary are mutually dependent. We developed a policy gradient method that exploits the value-based stopping structure and augments it with curriculum learning, gradually moving from forced continuation to adaptive stopping during training. Across a linear-Gaussian benchmark, a nonlinear non-conjugate test problem, and a convection-diffusion source detection problem, the method learned effective design–stopping policies. The experiments show that curriculum learning is most valuable when premature stopping prevents the policy from observing informative later-stage trajectories, as in the movement-constrained source detection problem with moderate experimental cost.
Several limitations motivate future work. First, evaluating terminal rewards at each decision stage can be expensive when posterior updates are costly or high dimensional. Second, alternative approaches that directly parameterize the stopping policy, without relying on the optimal stopping theorem (Theorem 3.1), may avoid circular dependencies and simplify training, albeit potentially at the expense of slower convergence or weaker theoretical guarantees. Finally, this paper focuses on information gain for parameter inference, but the same framework can accommodate goal-oriented utilities, such as improving predictions of future outcomes as in related work by Kleinegesse and Gutmann (2021). These extensions would broaden the applicability of optimal stopping to autonomous experimental systems with richer scientific and engineering objectives.
Overall, this work advances the development of autonomous experimental systems by enabling intelligent, resource-aware stopping decisions, improving the efficiency and robustness of sequential Bayesian experimental design.
@@ Appendix A Appendix
A.1 Proof of Theorem 3.1
Proof of Theorem 3.1.
@@
By induction, $V^{\pi,\psi}_{k}(s_{k})\geq V^{\pi,\psi^{\prime}}_{k}(s_{k})$ for all $k=0,\ldots,N-1$. In particular, $V^{\pi,\psi}_{0}(s_{0})\geq V^{\pi,\psi^{\prime}}_{0}(s_{0})$, which proves that the stopping policy in Theorem 3.1 is optimal. ∎
A.2 Proof of Theorem 3.2
Proof of Theorem 3.2.
We first prove the stopping sets are equivalent under two formulations. For convenience, we omit the notation of design policy $\pi$ and stopping policy $\psi$, but distinguish them using superscripts of incr and term respectively. The policy dependence in the expectations is also omitted for notational consistency. We first prove the following relationship between value functions:
@@
$V^{incr}_{k+1}(s_{k+1})=V^{term}_{k+1}(s_{k+1})-\left[D_{\mathrm{KL}}(p_{\theta|I_{k+1}}\,||\,p_{\theta|I_{0}})+\sum_{i=0}^{k}c_{i}(\xi_{i})\right].$ (41)
@@
Therefore, the stopping sets are equivalent in both formulations. In particular, the optimization objective is
@@
A.3 Proof of Theorem 3.3
To simplify notation, below we omit the subscript on $w$ and shorten $\mu_{k,w_{k}}(s_{k})$ to $\mu_{k,w}(s_{k})$, with the understanding that $w$ takes the same subscript as the $\mu$ function. We also suppress explicit conditioning in expectations whenever the underlying measure is clear from context. Throughout the proof, we impose the following regularity assumptions, which are standard in sequential decision problems with differentiable policies.
Assumption A.1.
(i) The observation noise $\varepsilon_{k}$ admits a density with respect to Lebesgue measure. (ii) The design policy $\mu_{k,w}(s_{k})$ is differentiable in $w$ for almost every $s_{k}$. (iii) The observation density $p(y_{k}\mid s_{k},\xi_{k})$ and the forward map $\mathcal{F}_{k}(s_{k},\xi_{k},y_{k})$ are differentiable in $\xi_{k}$. (iv) The derivatives in (iii) and the policy gradient $\nabla_{w}\mu_{k,w}$ are bounded.
Proof of Theorem 3.3.
The gradient of the expected utility can be expressed in terms of the value function:
@@
$V_{k}^{\pi_{w},\psi_{w}}(s_{k})$ $=\mathbf{1}_{s_{k}\in\mathcal{T}_{k,w}}r^{S}_{T}(s_{k})+(1-\mathbf{1}_{s_{k}\in\mathcal{T}_{k,w}})\mathbb{E}_{y_{k}}\left[r^{C}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k})+V^{\pi_{w},\psi_{w}}_{k+1}(\mathcal{F}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k}))\right]$
$=\mathbf{1}_{s_{k}\in\mathcal{T}_{k,w}}r^{S}_{T}(s_{k})+(1-\mathbf{1}_{s_{k}\in\mathcal{T}_{k,w}})Q_{k}(s_{k},w)$ (48)
where $Q_{k}(s_{k},w)=Q_{k}^{\pi_{w},\psi_{w}}(s_{k},\mu_{k,w}(s_{k}))=\mathbb{E}_{y_{k}}\left[r^{C}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k})+V^{\pi_{w},\psi_{w}}_{k+1}(\mathcal{F}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k}))\right]$, differentiating $V_{k}^{\pi_{w},\psi_{w}}(s_{k})$ with respect to $w$ yields
$\quad\nabla_{w}V^{\pi_{w},\psi_{w}}_{k}(s_{k})$ $=\nabla_{w}\mathbf{1}_{s_{k}\in\mathcal{T}_{k,w}}r^{S}_{T}(s_{k})-\nabla_{w}\mathbf{1}_{s_{k}\in\mathcal{T}_{k,w}}Q_{k}(s_{k},w)+(1-\mathbf{1}_{s_{k}\in\mathcal{T}_{k,w}})\nabla_{w}Q_{k}(s_{k},w)$
$=\nabla_{w}\mathbf{1}_{s_{k}\in\mathcal{T}_{k,w}}[r^{S}_{T}(s_{k})-Q_{k}(s_{k},w)]+\mathbf{1}_{s_{k}\notin\mathcal{T}_{k,w}}\nabla_{w}Q_{k}(s_{k},w).$ (49)
@@
$\partial\mathcal{T}_{k,w}=\{s_{k}\mid h(s_{k},w):=r^{S}_{T}(s_{k})-Q_{k}(s_{k},w)=0\},$ (50)
and correspondingly $\mathcal{T}_{k,w}=\{s_{k}\mid h(s_{k},w)>0\},\mathcal{T}^{c}_{k,w}=\{s_{k}\mid h(s_{k},w)\leq 0\}$. Since the indicator function changes value only on this boundary, its gradient can be represented in distributional form using a Dirac delta supported on $\partial\mathcal{T}_{k,w}$:
$\nabla_{w}\mathbf{1}_{s_{k}\in\mathcal{T}_{k,w}}=\delta(h(s_{k},w))\nabla_{w}h(s_{k},w).$ (51)
The first term in (49) is then reduced to $h(s_{k},w)\delta(h(s_{k},w))\nabla_{w}h(s_{k},w)$. Using the identity $x\delta(x)\equiv 0$, this boundary contribution vanishes. For the second term in (49),
$\nabla_{w}Q_{k}(s_{k},w)$ $=\nabla_{w}\mathbb{E}_{y_{k}}\left[r^{C}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k})+V^{\pi_{w},\psi_{w}}_{k+1}(\mathcal{F}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k}))\right]$
$=\nabla_{w}\mathbb{E}_{y_{k}}\left[r^{C}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k})\right]+\nabla_{w}\mathbb{E}_{y_{k}}\left[V^{\pi_{w},\psi_{w}}_{k+1}(\mathcal{F}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k}))\right],$ (52)
the latter term requires differentiating an expectation whose domain depends implicitly on $w$ through the stopping region. This can be handled using the Leibniz rule for parameter‑dependent integration domains, yielding both an interior gradient term and a boundary term proportional to the normal velocity of the moving boundary (see (Reed et al., 1972; Evans, 2022)).
$\quad\nabla_{w}\mathbb{E}_{y_{k}}\left[V^{\pi_{w},\psi_{w}}_{k+1}(\mathcal{F}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k}))\right]$
$=\nabla_{w}\int_{s_{k+1}}p(s_{k+1}|s_{k},\mu_{k,w}(s_{k}))V^{\pi_{w},\psi_{w}}_{k+1}(s_{k+1})\,ds_{k+1}$
$=\nabla_{w}\Bigg[\int_{\mathcal{T}_{k+1,w}}r_{T}^{S}(s_{k+1})p(s_{k+1}|s_{k},\mu_{k,w}(s_{k}))ds_{k+1}+\int_{\mathcal{T}^{c}_{k+1,w}}Q_{k+1}(s_{k+1},w)p(s_{k+1}|s_{k},\mu_{k,w}(s_{k}))ds_{k+1}\Bigg]$
$=\int_{\mathcal{T}_{k+1,w}}\nabla_{w}\left[r_{T}^{S}(s_{k+1})p(s_{k+1}|s_{k},\mu_{k,w}(s_{k}))\right]ds_{k+1}+\int_{\partial\mathcal{T}_{k+1,w}}r_{T}^{S}(s_{k+1})p(s_{k+1}|s_{k},\mu_{k,w}(s_{k}))\frac{\nabla_{w}h(s_{k+1},w)}{|\nabla_{s_{k+1}}h(s_{k+1},w)|}ds_{k+1}$
$\quad+\int_{\mathcal{T}^{c}_{k+1,w}}\nabla_{w}\left[Q_{k+1}(s_{k+1},w)p(s_{k+1}|s_{k},\mu_{k,w}(s_{k}))\right]ds_{k+1}$
$\quad-\int_{\partial\mathcal{T}_{k+1,w}}Q_{k+1}(s_{k+1},w)p(s_{k+1}|s_{k},\mu_{k,w}(s_{k}))\frac{\nabla_{w}h(s_{k+1},w)}{|\nabla_{s_{k+1}}h(s_{k+1},w)|}ds_{k+1}$
$=\int_{\mathcal{T}_{k+1,w}}\Bigg[p(s_{k+1}|s_{k},\mu_{k,w}(s_{k}))\nabla_{w}r_{T}^{S}(s_{k+1})+\nabla_{w}\mu_{k,w}(s_{k})\nabla_{\xi_{k}}p(s_{k+1}|s_{k},\xi_{k})\Big|_{\xi_{k}=\mu_{k,w}(s_{k})}r_{T}^{S}(s_{k+1})\Bigg]\,ds_{k+1}$
$\quad+\int_{\mathcal{T}^{c}_{k+1,w}}\Bigg[p(s_{k+1}|s_{k},\mu_{k,w}(s_{k}))\nabla_{w}Q_{k+1}(s_{k+1},w)+\nabla_{w}\mu_{k,w}(s_{k})\nabla_{\xi_{k}}p(s_{k+1}|s_{k},\xi_{k})\Big|_{\xi_{k}=\mu_{k,w}(s_{k})}Q_{k+1}(s_{k+1},w)\Bigg]\,ds_{k+1}$
$\quad+\int_{\partial\mathcal{T}_{k+1,w}}\left[r_{T}^{S}(s_{k+1})-Q_{k+1}(s_{k+1},w)\right]p(s_{k+1}|s_{k},\mu_{k,w}(s_{k}))\frac{\nabla_{w}h(s_{k+1},w)}{|\nabla_{s_{k+1}}h(s_{k+1},w)|}ds_{k+1}$
$=\int_{s_{k+1}}\Bigg[p(s_{k+1}|s_{k},\mu_{k,w}(s_{k}))\nabla_{w}V^{\pi_{w},\psi_{w}}_{k+1}(s_{k+1})+\nabla_{w}\mu_{k,w}(s_{k})\nabla_{\xi_{k}}p(s_{k+1}|s_{k},\xi_{k})\Big|_{\xi_{k}=\mu_{k,w}(s_{k})}V^{\pi_{w},\psi_{w}}_{k+1}(s_{k+1})\Bigg]\,ds_{k+1}$
$=\mathbb{E}_{y_{k}}\left[\nabla_{w}V^{\pi_{w},\psi_{w}}_{k+1}(\mathcal{F}_{k}(s_{k},\mu_{k,w}(s_{k}),y_{k}))\right],$ (53)
where the contribution of the boundary term vanishes since $r_{T}^{S}(s_{k+1})-Q_{k+1}(s_{k+1},w)=0$ on $\partial\mathcal{T}_{k+1,w}$. Combining the above identities yields
$\quad\nabla_{w}Q_{k}(s_{k},w)$